Subscribe to Standard Changes for IS Services and Systems feed
Announcements for OSU's Network and Computer Services
Updated: 51 min 8 sec ago

Data Center Maintenance in B210, A016 Kerr, and Milne

Tue, 01/20/2015 - 6:04am
Primary Contact: Joel Burks, We will be updating all three data centers starting at 5 am to 7 pm. There may be short network outages.

Emergency Maintenance of New Core Firewall

Fri, 01/02/2015 - 9:23am
Primary Contact: Joel Burks, Due to a bug in the current version of code on the new core firewall blades, an emergency upgrade was performed this morning at 7 am. There was no outage as redundancy measures were used.

Emergency Maintenance Exchange

Tue, 12/30/2014 - 12:38pm
Good afternoon, Tonight the IS Server Support team will performing emergency maintenance on one of the Exchange servers at 10PM to increase the service’s resources. We do not expect this maintenance to cause an outage. However some users may notice during the maintenance window that they are disconnected and reconnected as the mailboxes are migrated [...]

ClassroomSAN and DigitalAssets

Fri, 12/12/2014 - 12:56pm
Primary Contact: Mike Akey, On Wednesday December 17th, from 8am to 10am, the ClassroomSAN and will be unavailable while updates are installed and SAN maintenance is performed.


Fri, 12/05/2014 - 12:12am
Emergency maintenance complete. All clear. Typos From my Android phone

Emergency Maintenance SharePoint

Thu, 12/04/2014 - 2:38pm
Good afternoon, Tonight the IS Server Support team will performing emergency maintenance on one of the SharePoint servers at 10PM to address service issues. We expect this maintenance to cause an outage for about an hour while we add resources to the system. Who will this affect: All SharePoint users When will this happen: Friday [...]

LS Networks planned maintenance on 1 am to 4 am 12/04/2014

Mon, 12/01/2014 - 10:19am
Primary Contact: Joel Burks, The circuit provider for the Bend Hub has notified NERO of planned maintenance on Thursday, December 4th, between 0100-0400. Customers connected to the Bend Hub will experience intermittent connectivity during the maintenance window. Bend Hub Connected Customers: Central Oregon Community College High Desert ESD KLCC ODFW – Bend OSU Extension [...]

LS Networks planned maintenance on Midnight to 5 am 11/26/2014

Thu, 11/20/2014 - 9:44am
Primary Contact: Joel Burks, LSN fiber vendor will be performing an emergency maintenance on November 26th, 2014 00:01 to 02:00 PST in Astoria, OR. LSN fiber vendor will be re-splicing damaged fiber to avoid an potential outage.

11/11 morning maintenance complete

Tue, 11/11/2014 - 9:43am
Primary Contact: Tim Bateman, The maintenance on the core network from this morning is complete.

Maintenance on OB1

Tue, 11/11/2014 - 4:23am
Primary Contact: Tim Bateman, Beginning maintenance on OB1. There will be network connectivity loss in the core network this AM during maintenance.

migrating Aruba controller subnet to new core

Wed, 11/05/2014 - 9:22am
Primary Contact: Steven Heitmeyer, Aruba controller subnet has been successfully moved to the new core. I saw a loss of about 20 pings (= 20 seconds) while routing reconverged. One advantage is that now we have redundant routers. Also we will be able to upgrade to higher data rates in the future.

migrating Aruba controller subnet to new core

Wed, 11/05/2014 - 9:22am
Primary Contact: Steven Heitmeyer, I am beginning to migrate vlan 87 off of DRT1 (old core) and onto the Nexus 7K routers. There may be a short blip in wireless traffic during the transition.

2014-11-4 06:00-07:00

Wed, 11/05/2014 - 9:22am
Primary Contact: Richard Giesige, Network Operations will be conducting an upgrade to the VPN appliances to patch known security vulnerabilities. We do not expect any customer impact, but are sending an outage notice as their is a risk of failover issues during this upgrade period.

DCA Tools DB migration

Thu, 10/23/2014 - 4:59pm
Primary Contact: Mike Akey, We will be moving the backend database for the DCA web-based log searching tools during our regular maintenance window from 10pm to 12, Saturday Oct. 25th. During this time, the log searching via tools.nws will be unavailable.

Milne Commodity network outage

Fri, 10/17/2014 - 5:13pm
Primary Contact: Tim Bateman, Need to replace hardware on Milne building and floor switches to correct the building uplink throwing errors. Time: 8am-9am, Saturday, October 18th. Impact: Building commodity network for Milne will be down during part of the outage. Top floor network will be down the 2nd half of the outage.

Border Router Change 2014 October 14th 6 am to 7 am

Tue, 10/14/2014 - 6:06am
Primary Contact: Joel Burks, Performing border router maintenance. No outage expected.